OpenStudy (ajprincess):

Please help:) The iterative formula \(x_{i+1}=a_0+a_1x_1^2\) \((a_0, a_1\) positive ) is being used to solve the equation \(x=a_0+a_1x^2.\) What is the condition of convergence?

OpenStudy (anonymous):

did you try couchy?

OpenStudy (ajprincess):

sorry didn't get u?

OpenStudy (anonymous):

couchy convergence test: \[|x_{i+k}-x_{i}|<\epsilon\]

OpenStudy (ajprincess):

sorry i havnt learnt t.

OpenStudy (anonymous):

basicly it says that the far enough terms have distance less than any positive number

OpenStudy (perl):

i think here you assume that there is a limit

OpenStudy (perl):

one sec, let me work on it

OpenStudy (perl):

if limit exists, that means lim xn = L , then x_i+1 is roughly the same as x_i for large i

OpenStudy (perl):

so substitute L for xi and xi+1 L = a0 + a1 L ^2, it is a quadratic

OpenStudy (perl):

use quadratic formula , a1 L^2 - L + a0 = 0, where L = xi as i -> oo,

OpenStudy (perl):

L = [-(-1) + - sqrt ( 1 - 4*a1*a0)] / ( 2*a1)

OpenStudy (perl):

this has real solutions only when the discriminant is at least positive

OpenStudy (perl):

so the condition for convergence is , when a1*a0 <= 1/4

OpenStudy (perl):

also you have a typo in your question

OpenStudy (perl):

so for instance, a0 = 1/2 and a1 = 1/3 converges for any x

OpenStudy (perl):

for any initial seed x

OpenStudy (ajprincess):

Can u plz tell me what is the typo in the question? @perl

OpenStudy (perl):

your iterative formula

OpenStudy (perl):

it should say